OpenStudy (anonymous):

2x2 + 12x = 10 −3 ± square root 14 −3 ± 2square root 2 −3 ± square root 19 −3 ± square root 29

OpenStudy (mathlegend):

So I take it we must use the quadratic formula here..

OpenStudy (anonymous):

honestly I have no idea lol im lost as crap with this

OpenStudy (mathlegend):

\[2x ^{2}+12x-10=0\]

OpenStudy (mathlegend):

\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (mathlegend):

A = 2 B = 12 C = -10 Do you understand... how I chose those numbers first?

OpenStudy (anonymous):

not a clue man, I don't understand any of this stuff

OpenStudy (mathlegend):

\[ax ^{2}+bx+c = 2x ^{2}+12x-10\]

OpenStudy (mathlegend):

Now you see why 2 is a and 12 is b and -10 is c?

OpenStudy (anonymous):

uhhh..sure

OpenStudy (anonymous):

how do you find the answer

OpenStudy (mathlegend):

The formula... I am simply going to plug these numbers into their respective spots.

OpenStudy (mathlegend):

So it is imperative that you understand where I am pulling these numbers from... here we go...

OpenStudy (mathlegend):

\[\frac{ -12\pm \sqrt{12^{2}-4(2)(-10)} }{ 2(2) }\]

OpenStudy (mathlegend):

So now, if we simplify this... we should get...

OpenStudy (mathlegend):

\[\frac{ -12\pm \sqrt{224} }{ 4 }\]

OpenStudy (mathlegend):

\[\frac{ -3\pm 4\sqrt{14} }{ 4 }\]

OpenStudy (mathlegend):

The 4's reduce...

OpenStudy (mathlegend):

−3 ± square root 14 is your answer @tiffanyr96

OpenStudy (mathlegend):

Hopefully that helps somewhat... I understand that Math is everyone's cup of tea.